The sum of two consecutive even integers is three times four less than the smallest integer. What are the integers?
Answer by Eric Johnston(7)

Call the smallest integer x and then write it out as an equation thus:
x + (x+2) = -(3*4) + x
so 2x + 2 = -12 + x
so x = -14
The next even higher integer is -12
Check: sum = -14 -12 = -26
three times four less than lowest = - (3 * 4) - 14 = -26
Best regards, Eric.